OpenStudy (anonymous):

The initial population is 81. The population triples every 8 years. What is the yearly growth factor? What is the yearly growth rate?

OpenStudy (anonymous):

\[\sqrt[8]{3}\] which is about 1.1472 then the yearly growth rate is 114.72% or 14.72% increase every year

OpenStudy (anonymous):

how do you get the rate though? I don't understsnd that

OpenStudy (anonymous):

if the yearly growth rate is x, then every second year the population is (1+x) of last year. since the rate stays constant, we can say that : \[(1+x)^{8} = 3\] thus\[1+x=\sqrt[8]{3}\]